### v2

# Interplay between Electroweak Symmetry Breaking and Dark Matter Relic Density

Abstract
Models of Dark Matter must contend with the fact that the presence of electroweak symmetry breaking along the thermal evolution of the Universe modifies the masses, interactions and, thus, the thermally averaged cross sections. We study in detail the impact of taking (not taking) the presence of the electroweak symmetry breaking into account in the calculations of the Dark Matter relic density, providing a model-independent measure of such differences. By focusing on a particular model, we show that ignoring this effect can lead to the inclusion (exclusion) of wrong (viable) regions of parameter space.
preprint: IPPP/25/26
I Introduction
The spectrum of temperature fluctuations in the Cosmic Microwave Background has offered physicists an extraordinary opportunity to measure the relative abundances of the various energy components of the Universe. The structures that make up these fluctuations are consistent with the existence of a non-baryonic matter component that exceeds the amount of ordinary baryonic matter by a factor around five, offering, together with a number of experimental results from many other experimental sources [1], compelling evidence for the existence of Dark Matter (DM). The best determination of the present DM relic density is [2]
$\Omega h^{2}|_{\rm Obs}=0.1200\pm 0.0012,$ (1)
where $\Omega$ is the current energy density in the form of DM relative to the critical energy density of the Universe, and $h$ is today’s reduced Hubble constant in the units of 100100100\,100km/s/MPc.
Amongst the plethora of DM production mechanisms found in the literature [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], the thermal freeze-out scenario stands out for mirroring the thermal behavior of particles in the Standard Model (SM) [16]. [^1] Namely, photons and neutrinos. This mechanism assumes that in the early stages of our Universe, i.e. at very high temperatures, the interaction rate between DM and particles in the primordial thermal bath was strong enough to keep them in chemical and thermal equilibrium. However, as the Universe expanded and cooled, there arose a moment (the freeze-out instant) when the annihilation rate of DM into lighter particles became subdominant to the Hubble expansion rate, resulting in the decoupling of DM from the thermal plasma and the subsequent freeze-out of its comoving number density.
In the freeze-out scenario, the present relic density is computed by solving the Boltzmann Equation, which describes the change in the number density of the DM species over time. As an input, one needs to provide a cosmological history for our universe (contained in the Hubble parameter) and a set of thermally averaged DM depletion processes (coming from the particle physics model chosen, provided it has a DM candidate). The solution to the Boltzmann Equation can then be written implicitly as an integral over all time. With this information, one may predict the corresponding DM relic density and test it against the observed value in Eq. (1).
In practice, analytical and numerical solutions are hard to find, so one usually resorts to approximations to perform the aforementioned integral. One such approximation relies on the assumption that DM particles remain in thermal equilibrium with the primordial bath until the freeze-out instant, implying that the Boltzmann equation becomes dominated by DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale. In effect, this reduces the problem to integrating the Boltzmann Equation only from freeze-out to the present time [17, 18]. Since the relevant integration range lies at lower temperatures, the process becomes effectively low-energy, and finite temperature corrections are typically neglected in this scenario. This contrasts with other DM production mechanisms where these higher-order corrections play an important role in accurately predicting the present DM abundance [19, 20].
The standard approach, therefore, is to calculate all thermally averaged DM depletion processes in the low-energy regime and input them into the Boltzmann Equation. However, this reasoning misses the unavoidable fact that the Higgs boson acquires a non-zero vacuum expectation value (vev), triggering electroweak symmetry breaking (EWSB) at temperatures around $T_{\rm EWSB}\approx 160\,$GeV [21]. It is therefore expected that the particle masses and interaction rates change across the EWSB transition. In this sense, we propose an improved approach for calculating the DM relic density in which both before and after EWSB interactions are considered. As we will demonstrate, neglecting the effects of EWSB in the calculation of the DM relic density is especially concerning for DM candidates that freeze-out before EWSB and can lead to large discrepancies between the standard and improved approaches. We prove that the deviations from the standard approach are driven by a change in the freeze-out instant and/or by a change in the thermally averaged DM depletion processes. In addition, we provide an analytical way to quantify this deviation.
Time-varying masses have been studied before in the context of Variable-Mass Particles, where a scalar DM candidate is coupled to a quintessence field  [22, 23, 24]; in spontaneous freeze-out, where fermionic DM acquires its mass through an early phase transition, and freezes out due to its mass becoming larger than the plasma temperature [25]; in super-cool dark matter, where the universe undergoes a brief period of thermal inflation during which all particles become massless, leading to a dilution of the DM abundance [26]; and in the Flip-Flop vev mechanism, where an auxiliary scalar field acquires a temporary vacuum expectation value, inducing a period where DM is allowed to decay [27]. Furthermore, inclusion of thermally averaged DM depletion processes before and after EWSB has been considered in a model with a light scalar DM candidate produced via freeze-in [28], and in a two-component DM model where one particle freezes in and the other freezes out [29]. Yet, to the best of our knowledge, the effect has not been studied in the case of a single freeze-out DM component, and a systematic comparison between the standard and improved approaches and its physical implications is lacking in the literature.
The paper is organized as follows. In Section II, we start by discussing the standard way of computing thermal relic densities performed using the broken phase model throughout the thermal evolution of the Universe. Then, we repeat the calculation using the unbroken phase model before EWSB, and the broken phase model after EWSB, which we dub the improved approach. Figures of merit aimed at comparing the two approaches are introduced in Subsection II.3. In Section III, we introduce a simple model illustrating the differences between the two approaches, which are quantitatively explored in Section IV. The conclusions appear in Section V. Several detailed aspects of the calculation have been relegated to the appendices.
II Relic density: standard vs. improved approaches
II.1 Relic density: the standard approach
Consider a SM extension with a single DM candidate of mass $m_{\rm DM}$ which, at high enough temperatures, is kept in thermal and chemical equilibrium with the primordial bath through $2\to 2$ scattering processes. To calculate the present abundance of DM in our Universe, we make use of the Boltzmann Equation to track its number density $n_{\rm DM}$, over time. For practical purposes, it is useful to define the particle number density per comoving volume (or yield) as $Y_{\rm DM}\equiv n_{\rm DM}/s$, where $s$ is the entropy density of the Universe. The Boltzmann equation describing the time evolution of the yield in this scenario can be written as [17, 30]
$\frac{dY_{\rm DM}}{dx}=Z(x)\left(\overline{Y}(x)^{2}-Y_{\rm DM}^{2}\right),$ (2)
where $x\equiv m_{\rm DM}/T$ is the ratio of the DM mass to the temperature of the Universe,
$\overline{Y}(x)=\frac{45}{4\pi^{4}}\frac{x^{2}}{h_{\mathrm{eff}}(x)}K_{2}(x),$ (3)
is the equilibrium yield, and
$Z(x)=\sqrt{\frac{\pi}{45}}\frac{m_{\rm DM}M_{\rm Pl}}{x^{2}}\sqrt{g_{*}}\langle\sigma v\rangle(x).$ (4)
In Eqs. (3) and (4), the function $K_{2}(x)$ is the modified Bessel function of the second kind, $M_{\rm Pl}$ is the Planck mass, $\sqrt{g_{*}}=\frac{h_{\rm eff}}{\sqrt{g_{\rm eff}}}(1+\frac{T}{3h_{\rm eff}}\frac{dh_{\rm eff}}{dT})$, and $h_{\rm eff}$ and $g_{\rm eff}$ are the effective entropy and energy degrees of freedom, respectively. Further, $\langle\sigma v\rangle$ is the total thermal averaged DM annihilation cross-section.
Equation (2) can be solved numerically to find the present DM yield. However, to better grasp the main subject of this paper, it is useful to adopt the semi-analytical approach known as the freeze-out approximation [17, 30], where we assume a freeze-out instant $x_{f}\equiv m_{\rm DM}/T_{f}$, before which DM particles are in thermal equilibrium with the primordial bath (so $Y=\overline{Y}$), and after which they decouple, leading to their comoving number density freezing out (meaning $Y\gg\overline{Y}$). In this sense, it is convenient to reparameterize the yield as $Y_{\rm DM}\approx(1+\delta)\overline{Y}$, where $\delta$ is a monotonically increasing function of $x$. In this approximation, the decoupling instant $x_{f}$ can be numerically obtained by iteration through the following expression [17]
$x_{f}=\log\left(e^{x}\,\frac{\delta_{f}(2+\delta_{f})}{1+\delta_{f}}\frac{Z\,\overline{Y}}{\left(-\frac{\mathrm{d}\overline{Y}}{\mathrm{d}x}\right)}\right)_{x=x_{f}},$ (5)
where $\delta_{f}$ is a $\mathcal{O}(1)$ number [^2] For example, micrOMEGAs [31, 32] uses $\delta_{f}=1.5$. that dictates the amount by which the yield at freeze-out $Y_{f}$ differs from the equilibrium yield at the freeze-out instant $\overline{Y}(x_{f})$, through $Y_{f}=(1+\delta_{f})\overline{Y}(x_{f})$.
After the freeze-out instant, the standard approach is to neglect $\overline{Y}$ in Eq. (2) and integrate it explicitly for any $x>x_{f}$. In turn, the present-day DM yield is given by [17]:
$Y_{\rm DM}^{\rm today}\approx\frac{Y_{f}}{1+Y_{f}A_{x_{f}\to\infty}},$ (6)
where $A_{x_{1}\to x_{2}}$ is a function that quantifies the evolution of the yield between two points $x_{1}$ and $x_{2}$, and is given by
$A_{x_{1}\to x_{2}}=\int_{x_{1}}^{x_{2}}Z(x)\,\mathrm{d}x=\sqrt{\frac{\pi}{45}}m_{\mathrm{DM}}M_{\mathrm{Pl}}\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>}{x^{2}}\,\mathrm{d}x.$ (7)
Then, the present DM relic density is given by
$\Omega h^{2}\approx 2.744\times 10^{8}\,m_{\mathrm{DM}}\,Y_{\rm DM}^{\rm today}.$ (8)
II.2 Relic density: the improved approach
As explained in the previous section, the standard way to calculate the present DM relic density is to integrate the function $Z(x)$ from the $x$ value at decoupling ($x=x_{f}$) until the present epoch ($x\to\infty$). This is consistent with the idea that the Boltzmann equation is dominated by the DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale [18].
However, when EWSB occurs after the DM freeze-out, i.e., when $T_{f}>T_{\rm EWSB}$, special care must be taken in defining Eq. (7). This is because the integrand of Eq. (7) depends on both $m_{\rm DM}$ and $\langle\sigma v\rangle$, whose values change across the EWSB transition.
Note that EWSB is expected to occur at a temperature of approximately $T_{\rm EWSB}\sim 160\,\mathrm{GeV}$ [21], while the freeze-out of a typical scalar DM candidate takes place around $x_{f}\sim 25$ [17]. This implies that EWSB occurs after DM freeze-out when [29]
$x_{f}<x_{\rm EWSB}$$\implies$ $m_{\rm DM}>160\times 25\ {\rm GeV}\sim 4\,{\rm TeV}\,.$ (9)
One should therefore expect the standard approach to break down for DM masses above 4 TeV, motivating the need for an improved approach to more accurately calculate the present relic density for heavy DM candidates.
To properly define the improved approach — which incorporates the effects of EWSB — and to facilitate the comparison with the standard approach — where DM evolution is computed entirely in the electroweak broken phase, as is conventionally done in the dark matter literature —  we must introduce some notation. Let $\sigma\mathrm{v}\right>_{\rm bEWSB}$ be the thermal averaged cross-section before EWSB and $\sigma\mathrm{v}\right>_{\rm aEWSB}$ that after EWSB. Further, let $\widetilde{m}_{\rm DM}$ denote the DM mass in the electroweak unbroken phase and $m_{\rm DM}$ denote the DM mass in the electroweak broken phase. Then, the terminology we will use throughout this paper is
  1. (i)
    $\sigma v\right>_{S}$ : represents the total thermal averaged cross-section in the standard approach and is defined by the relation
    $\sigma v\right>_{S}(x)=\left<\sigma v\right>_{\rm aEWSB}(x),$ (10)
    throughout all cosmic history, i.e., over the full range of $x$.
  2. (ii)
    $\sigma v\right>_{I}$ : denotes the total thermal averaged cross-section in the improved prescription, which differentiates contributions from channels operating before and after EWSB, leading to the relation
[…]
    where $\Theta$ is the Heaviside theta function.
  3. (iii)
    $Y^{\infty}_{S}$ : denotes the present-day DM yield computed in the standard approach. Following Eqs. (6) and (7), it is given as:
    $Y^{\infty}_{S}=\frac{Y^{S}_{f}}{1+Y^{S}_{f}A^{S}_{{x^{S}_{f}}\to\infty}},$ (12)
    where the DM yield computed at the standard freeze-out instant $x_{f}^{S}$ is
    $Y_{f}^{S}=\left(1+\delta_{f}\right)\overline{Y}\left(x^{S}_{f}\right),$ (13)
    and
    $A^{S}_{{x_{1}}\to x_{2}}=\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,m_{\rm DM}\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>_{\rm aEWSB}}{x^{2}}\,\mathrm{d}x.$ (14)
  4. (iv)
    $Y^{\infty}_{I}$ : represents the present-day yield computed in the improved approach, and is defined by
    $Y^{\infty}_{I}=\frac{Y^{I}_{f}}{1+Y^{I}_{f}A^{I}_{{x^{I}_{f}}\to\infty}},$ (15)
    where the DM yield computed at the improved freeze-out instant $x_{f}^{I}$ is
    $Y_{f}^{I}=\left(1+\delta_{f}\right)\overline{Y}\left(x^{I}_{f}\right),$ (16)
    and
    $A^{I}_{{x_{1}}\to x_{2}}$ $=$ $\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,\widetilde{m}_{\mathrm{DM}}\,\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm bEWSB}}{x^{2}}\,\Theta(x_{\rm{EWSB}}-x)\,\mathrm{d}x$ (17)
    $+\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,m_{\mathrm{DM}}\,\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm aEWSB}}{x^{2}}\,\Theta(x-x_{\rm{EWSB}})\,\mathrm{d}x.$
  5. (v)
    $\Omega h^{2}|_{S}$ : is the standard DM relic density, defined as
    $\Omega h^{2}|_{S}=2.744\times 10^{8}\,m_{\mathrm{DM}}\,Y^{\infty}_{S}.$ (18)
  6. (vi)
    $\Omega h^{2}|_{I}$ : is the improved DM relic density, defined as
    $\Omega h^{2}|_{I}=2.744\times 10^{8}\,m_{\mathrm{DM}}\,Y^{\infty}_{I}.$ (19)
II.3 Comparing approaches
We use the relative difference between $\Omega h^{2}|_{S}$ and $\Omega h^{2}|_{I}$ to quantify the deviation between the standard and improved approaches. Based on the definitions of the relic densities introduced in Section II.2, we write the relative difference as
$\delta\Omega h^{2}\equiv\frac{\Omega h^{2}|_{I}-\Omega h^{2}|_{S}}{\Omega h^{2}|_{I}}=Y_{S}^{\infty}\left(\alpha-\beta\right),$ (20)
where we define
$\alpha$$\equiv$ $\frac{1}{Y^{S}_{f}}-\frac{1}{Y^{I}_{f}},$ (21)
$\beta$$\equiv$ $A^{I}_{{x^{I}_{f}}\to\infty}-A^{S}_{{x^{S}_{f}}\to\infty}.$ (22)
The sign of $\delta\Omega h^{2}$ is determined by the sign of $\left(\alpha-\beta\right)$. While $\alpha$ quantifies the difference in inverse of yields at freeze-out between the two approaches, $\beta$ is a function of the annihilation cross-sections, with the dependence following from Eqs. (14) and (17). Looking carefully at Eqs. (14) and (17), one may conclude that the quantity $A_{x_{\rm EWSB}\to\infty}$ is the same in both approaches. Further assuming both $\sigma v\right>$’s remain constant over temperature from the decoupling to EWSB, $\beta$ can be further simplified to [^3] In the illustrative model presented below, we have explicitly checked that Eq. (24) correctly predicts the sign of $\delta\Omega h^{2}$ for most (but not all) points in the parameter space. .
$\beta$ $=A^{I}_{{x^{I}_{f}}\to x_{\rm EWSB}}-A^{S}_{{x^{S}_{f}}\to x_{\rm EWSB}}$ (23)
[…]
Quantifying the relative error in this manner highlights the interplay between two significant factors: the annihilation cross-sections and the scalar masses, which together contribute to the shift in the final relic densities between the two approaches.
With these definitions established, in the following sections, we will analyze the evolution of the resulting dark matter relic densities and quantify the differences arising between the two approaches within a specific extension of the SM. To this end, we introduce next a simple model where the standard and improved approaches can have consequential differences. In any case, the possible influence of EWSB on the computation of DM relic density must be explored in the study of any freeze-out DM model, provided the mass of the DM candidate is above 4 TeV.
III A case study
III.1 Model definition
We extend the SM scalar sector by two real scalar singlets $\phi$ and $\chi$ [17, 30, 33]. The new scalars are stabilized by a $Z_{2}\times Z_{2}^{\prime}$ symmetry such that
$\chi\overset{Z_{2}}{\longrightarrow}-\chi\ \ \ \ \ \ \ \ {\rm and}\ \ \ \ \ \ \ \phi\overset{Z_{2}^{\prime}}{\longrightarrow}-\phi\,.$ (25)
Consequently, these fields couple not only to the SM Higgs doublet via portal couplings, but also to each other. Before EWSB, the most general renormalizable scalar potential reads:
$\begin{split}V_{\rm bEWSB}=&\ \mu_{H}^{2}H^{\dagger}H+\lambda_{H}\left(H^{\dagger}H\right)^{2}+\frac{1}{2}\mu_{\phi}^{2}\phi^{2}+\frac{1}{4!}\lambda_{\phi}\phi^{4}+\frac{1}{2}\mu_{\chi}^{2}\chi^{2}+\frac{1}{4!}\lambda_{\chi}\chi^{4}\\ &+\frac{1}{2}\lambda_{H\phi}\left(H^{\dagger}H\right)\phi^{2}+\frac{1}{2}\lambda_{H\chi}\left(H^{\dagger}H\right)\chi^{2}+\frac{1}{4}\lambda_{\phi\chi}\phi^{2}\chi^{2}.\end{split}$ (26)
At the minimum of the potential, the Hessian matrix is diagonal and the squared masses of the scalar fields come out as
$\widetilde{m}^{2}_{H}=\mu_{H}^{2}\,,\ \ \ \ \ \ \ \ \widetilde{m}^{2}_{\phi}=\mu_{\phi}^{2}\,,\ \ \ \ \ \ \ \ \widetilde{m}^{2}_{\chi}=\mu_{\chi}^{2}\,.$ (27)
After EWSB, both the Higgs doublet and $\phi$ acquire a vev, breaking the $SU(2)\times U(1)$ and $Z_{2}^{\prime}$ symmetries, respectively. In turn, the bilinear terms in $H$ and $\phi$ flip signs. Explicitly, the vacuum and field configuration after EWSB in the unitary gauge is
$H$ $\to\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v_{H}+h\end{pmatrix},$ $\phi$ $\to v_{\phi}+\phi\,,$ $\chi$ $\to\chi\,.$ (28)
Notice that this vacuum configuration allows for $h$ and $\phi$ to mix into the mass basis through
$\begin{pmatrix}h\\ \phi\end{pmatrix}=\begin{pmatrix}c_{\theta}&-s_{\theta}\\ s_{\theta}&c_{\theta}\end{pmatrix}\begin{pmatrix}h_{1}\\ h_{2}\end{pmatrix},$ (29)
where $c_{\theta}=\cos{\theta}$ and $s_{\theta}=\sin{\theta}$. We study this mixing in detail in App. A, and denote by $V_{\rm aEWSB}$ the potential after EWSB [^4] If $\lambda_{H\phi}=0$, then there is no coupling between the two sectors and $\theta=0$, c.f. Eq. (41).. The squared masses of the scalar fields read
$m_{h_{1}}^{2}$ $=$ $v_{H}^{2}\lambda_{H}+\frac{1}{6}v_{\phi}^{2}\lambda_{\phi}+\frac{1}{6}\sqrt{\left(6v_{H}^{2}\lambda_{H}-v_{\phi}^{2}\lambda_{\phi}\right)^{2}+\left(6v_{H}v_{\phi}\lambda_{H\phi}\right)^{2}},$ (30)
$m_{h_{2}}^{2}$ $=$ $v_{H}^{2}\lambda_{H}+\frac{1}{6}v_{\phi}^{2}\lambda_{\phi}-\frac{1}{6}\sqrt{\left(6v_{H}^{2}\lambda_{H}-v_{\phi}^{2}\lambda_{\phi}\right)^{2}+\left(6v_{H}v_{\phi}\lambda_{H\phi}\right)^{2}},$ (31)
$m_{\chi}^{2}$ $=$ $\mu_{\chi}^{2}+\frac{1}{2}v_{H}^{2}\lambda_{H\chi}+\frac{1}{2}v_{\phi}^{2}\lambda_{\phi\chi}.$ (32)
The scalars $h_{1}$ and $h_{2}$, besides coupling to $\chi$, have additional relevant interactions with $W$ and $Z$ gauge bosons, and SM fermions through
$\mathcal{L}\supset\left[\frac{\left(c_{\theta}h_{1}-s_{\theta}h_{2}\right)}{v_{H}}+\frac{\left(c_{\theta}h_{1}-s_{\theta}h_{2}\right)^{2}}{2v_{H}^{2}}\right]\left(2m_{W}^{2}W^{+}W^{-}+m_{Z}^{2}Z^{2}\right)-\left(c_{\theta}h_{1}-s_{\theta}h_{2}\right)\sum_{f}\frac{m_{f}}{v_{H}}\bar{f}f.$ (33)
We chose as input parameters for this model
$m_{h_{1}},\ m_{h_{2}},\ m_{\chi},\ v_{H},\ \sin\theta,\ \lambda_{\chi},\ \lambda_{H\phi},\ \lambda_{H\chi},\ \lambda_{\phi\chi},$ (34)
where we fix $v_{H}=246$ GeV, and identify $h_{1}$ with the $125$ GeV scalar found at the LHC [34, 35]. Note that the values of these parameters are chosen such that the scalar potential has a minimum both before and after EWSB. All other parameters can be written in terms of these; c.f. App. A.
III.2 Dark sector evolution
In this model, when $T>T_{\mathrm{EW}}$, both $\phi$ and $\chi$ are stable due to the unbroken $Z_{2}\times Z_{2}^{\prime}$ symmetry. This implies that there are two DM components, each forming its own dark sector. The possible annihilation channels are the quartic interactions $\phi\phi\to HH$, $\chi\chi\to HH$, and $\chi\chi\to\phi\phi$, shown in Fig. 1, which are controlled by the quartic couplings $\lambda_{H\phi}$, $\lambda_{H\chi}$ and $\lambda_{\phi\chi}$, respectively. Large values of these couplings keep the DM components in thermal equilibrium with the Higgs doublet $H$, which is itself in equilibrium with the SM bath because of sizable couplings with the SM gauge bosons and massless SM fermions.
Figure 1: Annihilation channels of scalar particles before EWSB
[…]
When $T<T_{\mathrm{EW}}$, EWSB kicks in and the scalars $H$ and $\phi$ acquire their respective vevs, the $Z_{2}^{\prime}$ symmetry is explicitly broken, no longer rendering $\phi$ as a DM component. In this electroweak broken phase, the mixing with the CP-even component of the Higgs doublet allows $h_{2}$ to decay into SM particles. These decay channels cause $Y_{h_{2}}$ to eventually drop to very low values as the temperature of the Universe begins to fall. This implies that, after EWSB, $\chi$ becomes the only DM component, still stabilized through the discrete symmetry assigned in Eq. (25). Also, because of the mixing between scalars in this regime, several additional channels open up as shown in Fig. 2, which feature the pair annihilation of $\chi$ into SM fermions and gauge bosons through $s$-channel processes mediated by $h_{1}$ and $h_{2}$.
Refer to caption   Refer to caption    Refer to caption   Refer to caption
Refer to caption    Refer to caption    Refer to caption
Figure 2: Dark matter annihilation channels after EWSB.
As introduced in Sec. II, the standard approach assumes DM to evolve entirely in the electroweak broken phase, while the improved approach distinguishes between the two phases — before EWSB and after EWSB — incorporating the full thermal history. It is worth reiterating here that the physical mass of DM is $m_{\chi}$ for the standard approach throughout the thermal history of the Universe, whereas the improved formalism considers DM mass as $\widetilde{m}_{\chi}$ until EWSB and $m_{\chi}$ afterward. Therefore, differences in relic density between the two approaches can arise either from changes in the DM annihilation cross section or from changes in the scalar masses across the two phases.
IV Results and observations
IV.1 Numerical scan
To better understand the extent to which the DM relic density differs between the standard and improved approaches, we performed a random scan of $10^{6}$ benchmark points to probe the region of parameter space bounded by:
$m_{\chi}$ $\in\left[1,100\right]\,\mathrm{TeV},$ $m_{h_{2}}$ $\in\left[250,1000\right]\,\mathrm{GeV},$
$\sin\theta$ $\in\left[-0.3,0.3\right],$ $\lambda_{\chi},\lambda_{H\phi},\lambda_{H\chi},\lambda_{\phi\chi}$$\in\pm\left[10^{-3},4\pi\right].$ (35)
As explained, the upper bound on the mass of $h_{2}$ is chosen to ensure it remains in thermal equilibrium with the SM bath until EWSB occurs. The range considered for the mixing angle $\theta$ is consistent with the latest experimental bounds on heavy Higgs-like scalar mixing [36]. Regarding quartic couplings, we scanned over their typical freeze-out values, ensuring they respect perturbativity and the boundedness from below conditions of App. B. We confirmed that, for values of $\lambda$’s below $10^{-3}$, there are no new features, but occasional numerical instabilities may occur.
To calculate the relic density of $\chi$ in the standard approach, we implemented, as usual, the model described in Sec. III in the electroweak broken phase. We used micrOMEGAs 6.0.5 [32] to numerically integrate the Boltzmann Equation and track the evolution of the yield of $\chi$ from $T=10^{5}$ GeV to $T=10^{-4}$ GeV. The standard relic density was then calculated using Eq. (18). On the other hand, the calculation of the relic density in the improved approach was performed in two steps. In the first step, we implemented the model before EWSB and used micrOMEGAs to integrate the Boltzmann Equation from $T=10^{5}$ GeV to $T=T_{\rm EWSB}$. In the second step, we used the model after EWSB and the previously calculated yield at $T=T_{\rm EWSB}$ as inputs to the darkOmegaNTR() function of micrOMEGAs, allowing us to integrate the Boltzmann Equation for $\chi$ from $T=T_{\rm EWSB}$ to $T=10^{-4}$ GeV. Finally, the improved relic density was then calculated using Eq. (19).
Figure 3: Relative deviation between approaches ($\delta\Omega h^{2}$) against $m_{\chi}$. Red points have $\Omega h^{2}|_{I}\approx 0.120$. Green points have $\Omega h^{2}|_{I}\approx 0.120$ and are allowed by Direct Detection constraints.
[…]
Masses Mixing and Couplings Relic Abundance
$m_{\chi}$ (TeV) $m_{\phi}$ (GeV) $\sin\theta$ $\lambda_{H\phi}$ $\lambda_{H\chi}$ $\lambda_{\phi\chi}$ $\Omega h^{2}|_{I}$ $\Omega h^{2}|_{S}$ $|\delta\Omega h^{2}|\,(\%)$
5.737 984 0.001 0.003 0.003 3.635 0.122721 0.125122 1.95
8.281 467 0.036 $-0.095$ $-0.248$ 5.103 0.118672 0.119259 0.49
[…]
Notice that the sign of $\delta\Omega h^{2}$ may either be positive or negative, as a consequence of the improved relic density lying above or below the standard relic density, respectively. The sign of the relative deviation between approaches can be predicted using Eq. (20). To confirm the extent to which the numerical evaluation performed in micrOMEGAs agrees with the improved semi-analytical formalism introduced in Sec. II.2, we compute $\alpha$ and $\beta$ using to Eqs (21) and (23) and use these values to position each point scanned on the $\alpha-\beta$ plane, as shown in Fig. 4. The color assigned to each point is based on the output from micrOMEGAs. We use red to indicate points with $\delta\Omega h^{2}>0$ and blue for points with $\delta\Omega h^{2}<0$. The black solid line delineates the boundary $\alpha=\beta$. The excellent agreement with the numerical results obtained from micrOMEGAs demonstrates that the improved approach reliably predicts the sign of the deviation.
Refer to captionRefer to caption
Figure 4: Sign of $\delta\Omega h^{2}$ plotted in the $\alpha-\beta$ plane.
To further highlight the need for a re-evaluation of DM relic density computations in TeV-scale dark sectors, in Fig. 5 we present the thermal history of $Y_{\chi}$ in two benchmark points. In these plots, the gray vertical line indicates the inverse of the EWSB temperature, and the gray horizontal line marks the yield corresponding to the observed relic density. The left panel illustrates a case where the final yield in the standard approach is overabundant and therefore excluded, but matches the observed relic density in the improved approach, thereby reopening otherwise disfavored regions of parameter space. The right panel features the opposite scenario, where a relic abundance deemed acceptable in the standard formalism turns out to be under-abundant in the improved approach, thereby constraining the parameter space further.
Figure 5: Left: A point allowed in the standard approach, but the improved approach forbids it. Right: A point discarded by the standard formalism but allowed in the improved approach.
IV.2 Picking a benchmark point
To better understand how changes in the dark matter annihilation cross-section and variations in scalar masses across the electroweak phase transition affect the relic density calculation, we consider a new benchmark scenario in which both effects are governed by a single parameter, $\lambda_{H\chi}$. We fix the remaining parameters as
$m_{\chi}=11.9\,\textrm{TeV}\,,\ \ m_{h_{2}}=895.1\,\textrm{GeV}\,,\ \ \sin{\theta}=0.11\,,\ \ \lambda_{\chi}=1.000\,,\ \ \lambda_{H\phi}=0.002\,,\ \ \lambda_{\phi\chi}=0.007\,.$ (36)
Figure 6: On the top (bottom) row of plots, the yield of $\chi$ (total thermal averaged cross-section) is plotted as a function of inverse temperature with $\lambda_{H\chi}$ increasing from left to right. The gray vertical line marks the inverse of $T_{\rm EWSB}$, and the grey horizontal line depicts the yield value corresponding to $\Omega h^{2}=0.12$.
In Fig. 6, we compare the DM yields predicted in the standard and improved approaches, as $\lambda_{H\chi}$ is increased. The top row of plots shows the evolution of the yield of $\chi$ as a function of $1/T$, with $\lambda_{H\chi}$ increasing from left to right. The black solid line represents the yield computed in the improved approach, the red dashed line represents the yield in the standard approach, the gray vertical line indicates the inverse of the EWSB temperature, and the gray horizontal line marks the yield corresponding to the observed relic density. For each value of $\lambda_{H\chi}$, we trace in the bottom row of plots the corresponding temperature evolution of $\langle\sigma v\rangle_{I}$ with a black solid line, and that of $\langle\sigma v\rangle_{S}$ with a red dashed line. It is interesting to note that $\chi$ always decouples from the SM bath earlier than in the improved approach, implying the relation $x_{f}^{S}\lesssim x_{f}^{I}$. Yet, regardless of the value of $\lambda_{H\chi}$, we always observe $\chi$ decoupling from the SM bath when $T>T_{\rm EWSB}$, as expected from the DM mass requirement in Eq. (9).
In the first column of Fig. 6, corresponding to $\lambda_{H\chi}=0.001$, we observe that $\chi$ freezes out with a smaller yield in the improved approach compared to the standard one, leading to an overall negative $\delta\Omega h^{2}$. This behavior can be understood as follows. At freeze-out, the standard and improved yields satisfy $Y_{f}^{S}>Y_{f}^{I}$, which implies a negative value for $\alpha$. On the other hand, before EWSB, we notice that $\langle\sigma v\rangle_{S}\lesssim\langle\sigma v\rangle_{I}$, which implies that $A^{S}_{{x^{I}_{f}}\to\infty}<A^{I}_{{x^{I}_{f}}\to\infty}$ and results in a positive $\beta$. For clarity, the particular values of $\alpha$ and $\beta$ for this benchmark point are marked by the white circular symbol in Fig. 4. A similar conclusion can be drawn for the other two benchmark points highlighted in Fig. 4, which correspond to the $\lambda_{H\chi}$ values used in the middle and right-most columns of Fig. 6. These two points, denoted by a triangle and a diamond in Fig. 4, exhibit positive $\alpha$. This arises from the fact that, before EWSB, the total thermal averaged cross-sections follow $\langle\sigma v\rangle_{S}>\langle\sigma v\rangle_{I}$, leading to a faster depletion of DM particles in the standard approach, and resulting in $Y_{f}^{S}<Y_{f}^{I}$. Due to the aforementioned tension between relevant parameters, $\beta$ remains negative in both cases. As a result, $\delta\Omega h^{2}$ turns out to be positive, in agreement with the values of $\delta\Omega h^{2}$ quoted on each row, and with the position of the solid black line (improved yield) lying above the red dashed one (standard yield). Note that for $\lambda_{H\chi}=4.9$, the improved relic density matches the observed value $\Omega h^{2}|_{\rm Obs}=0.12$ at the $3\sigma$ level, whereas the relic density computed in the standard approach does not. In fact, the standard relic density deviates from $\Omega h^{2}|_{\rm Obs}$ by $42\sigma$.
Figure 7: Contributions of individual annihilation channels to the improved total thermal averaged cross-section, as a function of inverse temperature.
For completeness, in Fig. 7 we also present the contributions of individual annihilation channels to the improved total thermal averaged cross-section, for each benchmark point studied in Fig. 6. As mentioned earlier, in the improved approach and before EWSB, the annihilation channels contributing to $Y_{\chi}$ are $\chi\chi\to\phi\phi$, and $\chi\chi\to HH$. When $\lambda_{\phi\chi}>\lambda_{H\chi}$, the total thermal averaged annihilation cross-section before EWSB is dominated by the process $\chi\chi\to\phi\phi$, which is in agreement with the first of this figure. However, as $\lambda_{H\chi}$ is increased, we observe the channel $\chi\chi\to HH$ becoming the dominant one.
V Conclusions
We propose a new approach to calculate the DM relic density in which the interplay between the occurrence of EWSB and the DM freeze-out instant is taken into account. In thermal freeze-out, one assumes that the DM particles remain in thermal equilibrium with the primordial bath until freeze-out occurs. In turn, this implies that the Boltzmann equation is solved by integrating from the freeze-out instant to the present time. In the standard approach, the DM particles and processes are always taken as those of the model’s broken phase. We have shown that in scenarios where DM freezes-out before EWSB, the use of the standard approach can lead to substantial discrepancies in the calculated value of the relic density. This has an important consequence for model building. By sticking to the standard approach, one can either exclude points in parameter space which are good, or permit points which in fact do not yield the correct relic abundance. In the limit, one may exclude a good model or permit a wrong one. The improved approach, described in our study, makes this distinction: we use the model in the unbroken phase, before EWSB; and the model in the broken phase, after EWSB. The differences between the two approaches have their origin in the thermally averaged cross section, which is modified by EWSB with a redefinition of masses and interactions. We were able to quantify the difference between the calculations in the two scenarios in a model independent way, in Eqs. (20)-(23).
We then made use of a simple model with two singlets and two independent symmetries; after EWSB, both the Higgs doublet and one of the scalars acquire vevs, breaking the $SU(2)\times U(1)$ and $Z_{2}^{\prime}$ symmetries, respectively. We chose to work with this model and vacuum configuration not only to simplify our analysis, but also because the differences between the standard and improved approaches were not that significant in the (simpler) real scalar singlet extension of the SM. Taking the DM mass above about 4 TeV after EWSB, we showed that indeed the two approaches lead to different relic densities. Moreover, this difference can be either negative or positive, implying the improved relic density can lie below or above the standard relic density, respectively.
It is important to note that these heavy DM particles are not excluded by experiment. The ongoing and planned direct and indirect detection experiments (see Ref. [1] for a comprehensive review) will most certainly give rise to an increase (or even exclusion) in the upper bound of the DM mass in this simple scenario with two singlets and two independent symmetries. However, it is easy to extend the model with, for example, an extra singlet, such that the couplings responsible for direct and indirect detection are not the same as the ones that determine the value of the relic density. We defer a detailed investigation of such extensions to future work.
The study also builds a strong case for testing TeV-scale phenomena in upcoming high-energy particle colliders such as the FCC-hh [40]. In fact, although models for which our approach is important have very heavy DM candidates, they could in principle be probed at future colliders. In this case the approach could prove instrumental in the reconstruction of a given potential from an extended scalar sector
VI Acknowledgments
This work is supported in part by the Portuguese Fundação para a Ciência e Tecnologia (FCT) through the PRR (Recovery and Resilience Plan), within the scope of the investment “RE-C06-i06 - Science Plus Capacity Building”, measure “RE-C06-i06.m02 - Reinforcement of financing for International Partnerships in Science, Technology and Innovation of the PRR”, under the projects with references 2024.01362.CERN and 2024.03328.CERN. The work of A.M. and J.P.S. is also supported by FCT under Contracts UIDB/00777/2020, and UIDP/00777/2020. The FCT projects are partially funded through POCTI (FEDER), COMPETE, QREN, and the EU. R.S. is partially supported by FCT under CFTC: UIDB/00618/2020 (https://doi.org/10.54499/UIDB/00618/2020), UIDP/00618/2020, (https://doi.org/10.54499/UIDP/00618/2020). A.M. is further supported by FCT with PhD Grant No. 2024.01340.BD. SC thanks Genevieve Belanger for useful discussions and acknowledges support from the UKRI Future Leader Fellowship DarkMAP.
Appendix A The potential after EWSB
The scalar potential in the unitary gauge after EWSB is given by
$V_{\rm aEWSB}$ $=$ $-\frac{1}{2}\mu_{H}^{2}(v_{H}+h)^{2}+\frac{1}{4}\lambda_{H}\left(v_{H}+h\right)^{2}-\frac{1}{2}\mu_{\phi}^{2}(v_{\phi}+\phi)^{2}+\frac{1}{4!}\lambda_{\phi}\phi^{4}$ (37)
$+\frac{1}{2}\mu_{\chi}^{2}\chi^{2}+\frac{1}{4!}\lambda_{\chi}\chi^{4}+\frac{1}{4}\lambda_{H\phi}(v_{H}+h)^{2}(v_{\phi}+\phi)^{2}$
$+\frac{1}{4}\lambda_{H\chi}(v_{H}+h)^{2}\chi^{2}+\frac{1}{4}\lambda_{\phi\chi}(v_{\phi}+\phi)^{2}\chi^{2}.$
The minimization of the potential after EWSB at the relevant extrema implies the relations
$\frac{\partial V_{\rm aEWSB}}{\partial h}=0$$\Leftrightarrow$ $\mu_{H}^{2}=v_{H}^{2}\lambda_{H}+\frac{1}{2}v_{\phi}^{2}\lambda_{H\phi},$ (38)
$\frac{\partial V_{\rm aEWSB}}{\partial\phi}=0$$\Leftrightarrow$ $\mu_{\phi}^{2}=\frac{1}{6}v_{\phi}^{2}\lambda_{\phi}+\frac{1}{2}v_{H}^{2}\lambda_{H\phi}.$ (39)
To study the scalar mass spectrum after EWSB, we collect the quadratic terms in the fields:
$V_{\rm aEWSB}\supset\frac{1}{2}\begin{pmatrix}h&\phi\end{pmatrix}\begin{pmatrix}2v_{H}^{2}\lambda_{H}&v_{H}v_{\phi}\lambda_{H\phi}\\ v_{H}v_{\phi}\lambda_{H\phi}&\frac{1}{3}v_{\phi}^{2}\lambda_{\phi}\end{pmatrix}\begin{pmatrix}h\\ \phi\end{pmatrix}+\frac{1}{2}\left(\mu_{\chi}^{2}+\frac{1}{2}v_{H}^{2}\lambda_{H\chi}+\frac{1}{2}v_{\phi}^{2}\lambda_{\phi\chi}\right)\chi^{2}.$ (40)
First, notice that this vacuum configuration allows $h$ and $\phi$ to mix. To rotate these fields to the physical eigenbasis, we introduce $\theta$ through Eq. (29). Then, the diagonalization of the squared mass matrix fixes the mixing angle at
$\tan\left(2\theta\right)=\frac{6\,v_{H}\,v_{\phi}\,\lambda_{H\phi}}{6v_{H}^{2}\lambda_{H}-v_{\phi}^{2}\lambda_{\phi}},$ (41)
and yields the masses in Eqs. (30)–(32). The external (input) parameters of this model are given in Eq. (34), while all the other parameters can be internally defined as
$v_{\phi}$ $=\frac{s_{\theta}c_{\theta}}{v_{H}\lambda_{H\phi}}\left(m_{h_{1}}^{2}-m_{h_{2}}^{2}\right),$ (42)
$\lambda_{H}$ $=\frac{1}{2v_{H}^{2}}\left(m_{h_{1}}^{2}c_{\theta}^{2}+m_{h_{2}}^{2}s_{\theta}^{2}\right),$ (43)
$\lambda_{\phi}$ $=\frac{3}{v_{\phi}^{2}}\left(m_{h_{1}}^{2}s_{\theta}^{2}+m_{h_{2}}^{2}c_{\theta}^{2}\right),$ (44)
$\mu_{H}^{2}$ $=v_{H}^{2}\lambda_{H}+\frac{1}{2}v_{\phi}^{2}\lambda_{H\phi},$ (45)
$\mu_{\phi}^{2}$ $=\frac{1}{6}v_{\phi}^{2}\lambda_{\phi}+\frac{1}{2}v_{H}^{2}\lambda_{H\phi},$ (46)
$\mu_{\chi}^{2}$ $=m_{\chi}^{2}-\frac{1}{2}v_{H}^{2}\lambda_{H\chi}-\frac{1}{2}v_{\phi}^{2}\lambda_{\phi\chi}.$ (47)
The scalar mass spectrum before EWSB is related to the mass spectrum after EWSB through Eq. (27). Explicitely these read
$2\,\widetilde{m}_{H}^{2}$ $=$ $\left(m_{h_{1}}^{2}c_{\theta}^{2}+m_{h_{2}}^{2}s_{\theta}^{2}\right)+\frac{s_{\theta}^{2}c_{\theta}^{2}}{v_{H}^{2}\lambda_{H\phi}}\left(m_{h_{1}}^{2}-m_{h_{2}}^{2}\right)^{2},$ (48)
$2\,\widetilde{m}_{\phi}^{2}$ $=$ $\left(m_{h_{1}}^{2}s_{\theta}^{2}+m_{h_{2}}^{2}c_{\theta}^{2}\right)+v_{H}^{2}\lambda_{H\phi},$ (49)
$\widetilde{m}_{\chi}^{2}$ $=$ $m_{\chi}^{2}-\frac{1}{2}v_{H}^{2}\lambda_{H\chi}-\frac{1}{2}\frac{\lambda_{\phi\chi}}{\lambda_{H\phi}^{2}}\frac{s_{\theta}^{2}c_{\theta}^{2}}{v_{H}^{2}}\left(m_{h_{1}}^{2}-m_{h_{2}}^{2}\right)^{2}.$ (50)
Notice that the values of the external parameters must be such that the scalar potential has a minimum both before and after EWSB. This condition is met when, before EWSB, the Hessian matrix is semi-definite, i.e. when $\widetilde{m}_{H}^{2}$, $\widetilde{m}_{\phi}^{2}$ and $\widetilde{m}_{\chi}^{2}$ are positive.
Appendix B Boundedness from below
For the scalar potential studied in this paper to be bounded from below, we must ensure that [41, 42]:
$\lambda_{H}>0\,,\quad\quad\lambda_{\phi}>0\,,\quad\quad\lambda_{\chi}>0\,.$ (51)
$\lambda_{H\chi}+\sqrt{\frac{2}{3}\lambda_{H}\lambda_{\chi}}>0\,,\quad\quad\lambda_{H\phi}+\sqrt{\frac{2}{3}\lambda_{H}\lambda_{\phi}}>0\,,\quad\quad 3\lambda_{\phi\chi}+\sqrt{\lambda_{\phi}\lambda_{\chi}}>0\,.$ (52)
Then, when $3\lambda_{\phi\chi}<\sqrt{\lambda_{\chi}\lambda_{\phi}}$ it is sufficient to require
$\lambda_{H\phi}<-\lambda_{\phi\chi}\sqrt{\frac{6\lambda_{H}}{\lambda_{\chi}}}\,,\quad\quad\quad 9\lambda_{H\phi}\lambda_{\phi\chi}-3\lambda_{H\chi}\lambda_{\chi}<\sqrt{3\left(3\lambda_{H\phi}^{2}-2\lambda_{H}\lambda_{\phi}\right)\left(9\lambda_{\phi\chi}^{2}-\lambda_{\phi}\lambda_{\chi}\right)},$ (53)
Appendix C Direct Detection
The DM-nucleon spin-independent elastic scattering cross-section ($\sigma_{\mathrm{SI}}$), in this model reads [43]
$\sigma_{\mathrm{SI}}$ $=$ $\frac{1}{\pi}\left(\frac{m_{N}^{2}}{m_{N}+m_{\chi}}\right)^{2}\frac{f_{N}^{2}}{v_{H}^{2}}\left|\frac{\cos\theta\lambda_{h_{1}\chi\chi}}{m_{h_{1}}^{2}}-\frac{\sin\theta\lambda_{h_{2}\chi\chi}}{m_{h_{2}}^{2}}\right|^{2},$ (54)
where $f_{N}\approx 0.308$ is the Higgs-nucleon effective coupling [44], $m_{N}\approx 0.938$ GeV is the nucleon mass, and
$\lambda_{h_{1}\chi\chi}$ $=$ $\cos\theta\,v_{H}\lambda_{H\chi}+\sin\theta\,v_{\phi}\lambda_{\phi\chi},$ (55)
$\lambda_{h_{2}\chi\chi}$ $=$ $-\sin\theta\,v_{H}\lambda_{H\chi}+\cos\theta\,v_{\phi}\lambda_{\phi\chi}.$ (56)

### v3

@@
 The spectrum of temperature fluctuations in the Cosmic Microwave Background has offered physicists an extraordinary opportunity to measure the relative abundances of the various energy components of the Universe.
-The structures that make up these fluctuations are consistent with the existence of a non-baryonic matter component that exceeds the amount of ordinary baryonic matter by a factor around five, offering, together with a number of experimental results from many other experimental sources [1], compelling evidence for the existence of Dark Matter (DM).
-The best determination of the present DM relic density is [2]
+The structures that make up these fluctuations are consistent with the existence of a non-baryonic matter component that exceeds the amount of ordinary baryonic matter by a factor around five, offering, together with a number of experimental results from many other experimental sources Cirelli:2024ssz , compelling evidence for the existence of Dark Matter (DM).
+The best determination of the present DM relic density is Planck:2018vyg
 $\Omega h^{2}|_{\rm Obs}=0.1200\pm 0.0012,$ (1)
-where $\Omega$ is the current energy density in the form of DM relative to the critical energy density of the Universe, and $h$ is today’s reduced Hubble constant in the units of 100100100\,100km/s/MPc.
-Amongst the plethora of DM production mechanisms found in the literature [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], the thermal freeze-out scenario stands out for mirroring the thermal behavior of particles in the Standard Model (SM) [16].
+where $\Omega$ is the current energy density in the form of DM relative to the critical energy density of the Universe, and $h$ is today’s reduced Hubble constant in the units of 100100\,km/s/MPc.
+Amongst the plethora of DM production mechanisms found in the literature DEramo:2010keq ; Hochberg:2014dra ; Kuflik:2015isi ; DAgnolo:2015ujb ; Pappadopulo:2016pkp ; Garny:2017rxs ; DAgnolo:2020mpt ; Smirnov:2020zwf ; Fitzpatrick:2020vba ; Kramer:2020sbb ; Hryczuk:2021qtz ; Bringmann:2021tjr , the thermal freeze-out scenario stands out for mirroring the thermal behavior of particles in the Standard Model (SM) Bauer:2017qwy .
 [^1] Namely, photons and neutrinos.
@@
 One such approximation relies on the assumption that DM particles remain in thermal equilibrium with the primordial bath until the freeze-out instant, implying that the Boltzmann equation becomes dominated by DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale.
-In effect, this reduces the problem to integrating the Boltzmann Equation only from freeze-out to the present time [17, 18].
+In effect, this reduces the problem to integrating the Boltzmann Equation only from freeze-out to the present time Cline:2013gha ; Arcadi:2024ukq .
 Since the relevant integration range lies at lower temperatures, the process becomes effectively low-energy, and finite temperature corrections are typically neglected in this scenario.
-This contrasts with other DM production mechanisms where these higher-order corrections play an important role in accurately predicting the present DM abundance [19, 20].
+This contrasts with other DM production mechanisms where these higher-order corrections play an important role in accurately predicting the present DM abundance Baker:2017zwx ; Chakrabarty:2022bcn .
 The standard approach, therefore, is to calculate all thermally averaged DM depletion processes in the low-energy regime and input them into the Boltzmann Equation.
-However, this reasoning misses the unavoidable fact that the Higgs boson acquires a non-zero vacuum expectation value (vev), triggering electroweak symmetry breaking (EWSB) at temperatures around $T_{\rm EWSB}\approx 160\,$GeV [21].
+However, this reasoning misses the unavoidable fact that the Higgs boson acquires a non-zero vacuum expectation value (vev), triggering electroweak symmetry breaking (EWSB) at temperatures around $T_{\rm EWSB}\approx 160\,$GeV Laine:2015kra .
 It is therefore expected that the particle masses and interaction rates change across the EWSB transition.
 In this sense, we propose an improved approach for calculating the DM relic density in which both before and after EWSB interactions are considered.
-As we will demonstrate, neglecting the effects of EWSB in the calculation of the DM relic density is especially concerning for DM candidates that freeze-out before EWSB and can lead to large discrepancies between the standard and improved approaches.
+As we will demonstrate, neglecting the effects of EWSB in the calculation of the DM relic density is especially concerning for DM candidates that freeze-out before EWSB, i.e.
+for DM candidates with a mass above $\sim 4$ TeV, and can lead to large discrepancies between the standard and improved approaches.
 We prove that the deviations from the standard approach are driven by a change in the freeze-out instant and/or by a change in the thermally averaged DM depletion processes.
 In addition, we provide an analytical way to quantify this deviation.
-Time-varying masses have been studied before in the context of Variable-Mass Particles, where a scalar DM candidate is coupled to a quintessence field  [22, 23, 24]; in spontaneous freeze-out, where fermionic DM acquires its mass through an early phase transition, and freezes out due to its mass becoming larger than the plasma temperature [25]; in super-cool dark matter, where the universe undergoes a brief period of thermal inflation during which all particles become massless, leading to a dilution of the DM abundance [26]; and in the Flip-Flop vev mechanism, where an auxiliary scalar field acquires a temporary vacuum expectation value, inducing a period where DM is allowed to decay [27].
-Furthermore, inclusion of thermally averaged DM depletion processes before and after EWSB has been considered in a model with a light scalar DM candidate produced via freeze-in [28], and in a two-component DM model where one particle freezes in and the other freezes out [29].
-Yet, to the best of our knowledge, the effect has not been studied in the case of a single freeze-out DM component, and a systematic comparison between the standard and improved approaches and its physical implications is lacking in the literature.
+Time-varying masses have been studied before in the context of Variable-Mass Particles, where a scalar DM candidate is coupled to a quintessence field  Anderson:1997un ; Rosenfeld:2005pw ; Franca:2003zg ; in spontaneous freeze-out, where fermionic DM acquires its mass through an early phase transition, and freezes out due to its mass becoming larger than the plasma temperature Heurtier:2019beu ; in super-cool dark matter, where the universe undergoes a brief period of thermal inflation during which all particles become massless, leading to a dilution of the DM abundance Hambye:2018qjv ; and in the Flip-Flop vev mechanism, where an auxiliary scalar field acquires a temporary vacuum expectation value, inducing a period where DM is allowed to decay Baker:2018vos .
+Furthermore, inclusion of thermally averaged DM depletion processes before and after EWSB has been considered in a model with a light scalar DM candidate produced via freeze-in Heeba:2018wtf , and in a two-component DM model where one particle freezes in and the other freezes out Bhattacharya:2021rwh .
+However, to the best of our knowledge, a systematic comparison between the standard and improved approaches and its physical implications is lacking in the literature.
 The paper is organized as follows.
@@
 For practical purposes, it is useful to define the particle number density per comoving volume (or yield) as $Y_{\rm DM}\equiv n_{\rm DM}/s$, where $s$ is the entropy density of the Universe.
-The Boltzmann equation describing the time evolution of the yield in this scenario can be written as [17, 30]
-$\frac{dY_{\rm DM}}{dx}=Z(x)\left(\overline{Y}(x)^{2}-Y_{\rm DM}^{2}\right),$ (2)
-where $x\equiv m_{\rm DM}/T$ is the ratio of the DM mass to the temperature of the Universe,
-$\overline{Y}(x)=\frac{45}{4\pi^{4}}\frac{x^{2}}{h_{\mathrm{eff}}(x)}K_{2}(x),$ (3)
+The Boltzmann equation describing the evolution of the yield in this scenario can be written as Cline:2013gha ; Steigman:2012nb
+$\frac{dY_{\rm DM}}{dy}=Z(y)\left(\overline{Y}(y)^{2}-Y_{\rm DM}^{2}\right),$ (2)
+where [^2] Usually, the independent variable is defined as the ratio of the DM mass to the temperature of the Universe, i.e.
+$x=m_{\rm DM}/T$.
+However, since we consider a scenario where $m_{\rm DM}$ changes discontinuously across a phase transition, this definition is ill-defined, and we work with $y=1/T$ instead.
+$y\equiv 1/T$ is the ratio of the DM mass to the temperature of the Universe,
+$\overline{Y}(y)\equiv\frac{45}{4\pi^{4}}\frac{m_{\rm DM}^{2}y^{2}}{h_{\mathrm{eff}}(y)}K_{2}\left(m_{\rm DM}\,y\right),$ (3)
 is the equilibrium yield, and
-$Z(x)=\sqrt{\frac{\pi}{45}}\frac{m_{\rm DM}M_{\rm Pl}}{x^{2}}\sqrt{g_{*}}\langle\sigma v\rangle(x).$ (4)
+$Z(y)\equiv\sqrt{\frac{\pi}{45}}M_{\rm Pl}\frac{\sqrt{g_{*}}\,\langle\sigma v\rangle}{m_{\rm DM}\,y^{2}}.$ (4)
 In Eqs.
-(3) and (4), the function $K_{2}(x)$ is the modified Bessel function of the second kind, $M_{\rm Pl}$ is the Planck mass, $\sqrt{g_{*}}=\frac{h_{\rm eff}}{\sqrt{g_{\rm eff}}}(1+\frac{T}{3h_{\rm eff}}\frac{dh_{\rm eff}}{dT})$, and $h_{\rm eff}$ and $g_{\rm eff}$ are the effective entropy and energy degrees of freedom, respectively.
-Further, $\langle\sigma v\rangle$ is the total thermal averaged DM annihilation cross-section.
+(3) and (4), the special function $K_{2}$ is the modified Bessel function of the second kind, $M_{\rm Pl}$ is the Planck mass, $\sqrt{g_{*}}=\frac{h_{\rm eff}}{\sqrt{g_{\rm eff}}}(1+\frac{T}{3h_{\rm eff}}\frac{dh_{\rm eff}}{dT})$, and $h_{\rm eff}$ and $g_{\rm eff}$ are the effective entropy and energy degrees of freedom, respectively.
+Further, $\langle\sigma v\rangle$ is the total thermal-averaged DM annihilation cross-section, computed through the Gondolo-Gelmini formula Gondolo:1990dk :
+$\langle\sigma v\rangle\equiv\frac{y}{8m_{\rm DM}^{4}K_{2}^{2}\left(m_{\rm DM}\,y\right)}\int_{4m_{\rm DM}^{2}}^{\infty}\sigma(s-4m_{\rm DM}^{2})\sqrt{s}K_{1}\left(\sqrt{s}\,y\right)\mathrm{d}s.$ (5)
+Here, the special function $K_{1}$ is the modified Bessel function of the first kind, and $\sigma$ is the total DM annihilation scattering cross-section.
 Equation (2) can be solved numerically to find the present DM yield.
-However, to better grasp the main subject of this paper, it is useful to adopt the semi-analytical approach known as the freeze-out approximation [17, 30], where we assume a freeze-out instant $x_{f}\equiv m_{\rm DM}/T_{f}$, before which DM particles are in thermal equilibrium with the primordial bath (so $Y=\overline{Y}$), and after which they decouple, leading to their comoving number density freezing out (meaning $Y\gg\overline{Y}$).
-In this sense, it is convenient to reparameterize the yield as $Y_{\rm DM}\approx(1+\delta)\overline{Y}$, where $\delta$ is a monotonically increasing function of $x$.
-In this approximation, the decoupling instant $x_{f}$ can be numerically obtained by iteration through the following expression [17]
-$x_{f}=\log\left(e^{x}\,\frac{\delta_{f}(2+\delta_{f})}{1+\delta_{f}}\frac{Z\,\overline{Y}}{\left(-\frac{\mathrm{d}\overline{Y}}{\mathrm{d}x}\right)}\right)_{x=x_{f}},$ (5)
-where $\delta_{f}$ is a $\mathcal{O}(1)$ number [^2] For example, micrOMEGAs [31, 32] uses $\delta_{f}=1.5$.
-that dictates the amount by which the yield at freeze-out $Y_{f}$ differs from the equilibrium yield at the freeze-out instant $\overline{Y}(x_{f})$, through $Y_{f}=(1+\delta_{f})\overline{Y}(x_{f})$.
-After the freeze-out instant, the standard approach is to neglect $\overline{Y}$ in Eq. (2) and integrate it explicitly for any $x>x_{f}$. In turn, the present-day DM yield is given by [17]:
-$Y_{\rm DM}^{\rm today}\approx\frac{Y_{f}}{1+Y_{f}A_{x_{f}\to\infty}},$ (6)
-where $A_{x_{1}\to x_{2}}$ is a function that quantifies the evolution of the yield between two points $x_{1}$ and $x_{2}$, and is given by
-$A_{x_{1}\to x_{2}}=\int_{x_{1}}^{x_{2}}Z(x)\,\mathrm{d}x=\sqrt{\frac{\pi}{45}}m_{\mathrm{DM}}M_{\mathrm{Pl}}\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>}{x^{2}}\,\mathrm{d}x.$ (7)
+All the results and parameter space presented in our study are based on solving the Boltzmann equation numerically, adapted to include the phase transition effects.
+However, to better grasp the main subject of this paper, it is useful to recall the semi-analytical freeze-out approximation Cline:2013gha ; Steigman:2012nb , which we employ only to demonstrate the subtleties of the analysis.
+In this approximation, one assumes a freeze-out instant $y_{f}\equiv 1/T_{f}$, before which before which DM particles are in thermal equilibrium with the primordial bath (so $Y=\overline{Y}$), and after which they decouple, leading to their comoving number density freezing out (meaning $Y\gg\overline{Y}$).
+In this sense, it is convenient to reparameterize the yield as $Y_{\rm DM}\approx(1+\delta)\overline{Y}$, where $\delta$ is a monotonically increasing function of $y$.
+In this approximation, the decoupling instant $y_{f}$ can be numerically obtained by iteration through the following expression Cline:2013gha
+$y_{f}=\log\left(e^{y}\,\frac{\delta_{f}(2+\delta_{f})}{1+\delta_{f}}\frac{Z\,\overline{Y}}{\left(-\frac{\mathrm{d}\overline{Y}}{\mathrm{d}y}\right)}\right)_{y=y_{f}},$ (6)
+where $\delta_{f}$ is a $\mathcal{O}(1)$ number [^3] For example, micrOMEGAs Belanger:2001fz ; Alguero:2023zol uses $\delta_{f}=1.5$.
+that dictates the amount by which the yield at freeze-out $Y_{f}$ differs from the equilibrium yield at the freeze-out instant $\overline{Y}(y_{f})$, through $Y_{f}=(1+\delta_{f})\overline{Y}(y_{f})$.
+After the freeze-out instant, the standard approach is to neglect $\overline{Y}$ in Eq. (2) and integrate it explicitly for all $y>y_{f}$. In turn, the present-day DM yield is given by Cline:2013gha :
+$Y_{\rm DM}^{\rm today}\approx\frac{Y_{f}}{1+Y_{f}A_{y_{f}\to\infty}},$ (7)
+where $A_{y_{1}\to y_{2}}$ is a function that quantifies the evolution of the yield between two points $y_{1}$ and $y_{2}$, and is given by
+$A_{y_{1}\to y_{2}}=\int_{y_{1}}^{y_{2}}Z(y)\,\mathrm{d}y=\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>}{m_{\mathrm{DM}}\,y^{2}}\,\mathrm{d}y.$ (8)
 Then, the present DM relic density is given by
@@
 II.2 Relic density: the improved approach
-As explained in the previous section, the standard way to calculate the present DM relic density is to integrate the function $Z(x)$ from the $x$ value at decoupling ($x=x_{f}$) until the present epoch ($x\to\infty$).
-This is consistent with the idea that the Boltzmann equation is dominated by the DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale [18].
+As explained in the previous section, the standard way to calculate the present DM relic density is to integrate the function $Z(y)$ from the $y$ value at decoupling ($y=y_{f}$) until the present epoch ($y\to\infty$).
+This is consistent with the idea that the Boltzmann equation is dominated by the DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale Arcadi:2024ukq .
 However, when EWSB occurs after the DM freeze-out, i.e., when $T_{f}>T_{\rm EWSB}$, special care must be taken in defining Eq.
@@
 (8) depends on both $m_{\rm DM}$ and $\langle\sigma v\rangle$, whose values change across the EWSB transition.
-Note that EWSB is expected to occur at a temperature of approximately $T_{\rm EWSB}\sim 160\,\mathrm{GeV}$ [21], while the freeze-out of a typical scalar DM candidate takes place around $x_{f}\sim 25$ [17].
-This implies that EWSB occurs after DM freeze-out when [29]
-$x_{f}<x_{\rm EWSB}$$\implies$ $m_{\rm DM}>160\times 25\ {\rm GeV}\sim 4\,{\rm TeV}\,.$ (9)
+Note that EWSB is expected to occur at a temperature of approximately $T_{\rm EWSB}\sim 160\,\mathrm{GeV}$ Laine:2015kra , while the freeze-out of a typical scalar DM candidate takes place around $T_{f}\sim m_{\mathrm{DM}}/25$ Cline:2013gha .
+This implies that EWSB occurs after DM freeze-out when Bhattacharya:2021rwh
+$T_{f}>T_{\rm EWSB}\Leftrightarrow y_{f}<y_{\rm EWSB}$ $\implies$ $m_{\rm DM}>160\times 25\ {\rm GeV}\sim 4\,{\rm TeV}\,.$ (10)
 One should therefore expect the standard approach to break down for DM masses above 4 TeV, motivating the need for an improved approach to more accurately calculate the present relic density for heavy DM candidates.
 To properly define the improved approach — which incorporates the effects of EWSB — and to facilitate the comparison with the standard approach — where DM evolution is computed entirely in the electroweak broken phase, as is conventionally done in the dark matter literature —  we must introduce some notation.
-Let $\sigma\mathrm{v}\right>_{\rm bEWSB}$ be the thermal averaged cross-section before EWSB and $\sigma\mathrm{v}\right>_{\rm aEWSB}$ that after EWSB.
+Let $\left<\sigma\mathrm{v}\right>_{\rm bEWSB}$ be the thermal averaged cross-section before EWSB and $\left<\sigma\mathrm{v}\right>_{\rm aEWSB}$ that after EWSB.
 Further, let $\widetilde{m}_{\rm DM}$ denote the DM mass in the electroweak unbroken phase and $m_{\rm DM}$ denote the DM mass in the electroweak broken phase.
@@
   1. (i)
-    $\sigma v\right>_{S}$ : represents the total thermal averaged cross-section in the standard approach and is defined by the relation
-    $\sigma v\right>_{S}(x)=\left<\sigma v\right>_{\rm aEWSB}(x),$ (10)
-    throughout all cosmic history, i.e., over the full range of $x$.
+    $\left<\sigma v\right>_{S}$ : represents the total thermal averaged cross-section in the standard approach and is defined by the relation
+    $\left<\sigma v\right>_{S}(y)=\left<\sigma v\right>_{\rm aEWSB}(y),$ (11)
+    throughout all cosmic history, i.e., over the full range of $y$.
   2. (ii)
-    $\sigma v\right>_{I}$ : denotes the total thermal averaged cross-section in the improved prescription, which differentiates contributions from channels operating before and after EWSB, leading to the relation
-[…]
+    $\left<\sigma v\right>_{I}$ : denotes the total thermal averaged cross-section in the improved prescription, which differentiates contributions from channels operating before and after EWSB, leading to the relation
+    $\left<\sigma v\right>_{I}(y)=\left<\sigma v\right>_{\rm bEWSB}(y)\,\Theta(y_{\rm{EWSB}}-y)+\left<\sigma v\right>_{\rm aEWSB}(y)\,\Theta(y-y_{\rm{EWSB}}),$ (12)
     where $\Theta$ is the Heaviside theta function.
@@
     $Y^{\infty}_{S}$ : denotes the present-day DM yield computed in the standard approach. Following Eqs. (7) and (8), it is given as:
-    $Y^{\infty}_{S}=\frac{Y^{S}_{f}}{1+Y^{S}_{f}A^{S}_{{x^{S}_{f}}\to\infty}},$ (12)
-    where the DM yield computed at the standard freeze-out instant $x_{f}^{S}$ is
-    $Y_{f}^{S}=\left(1+\delta_{f}\right)\overline{Y}\left(x^{S}_{f}\right),$ (13)
+    $Y^{\infty}_{S}=\frac{Y^{S}_{f}}{1+Y^{S}_{f}A^{S}_{{y^{S}_{f}}\to\infty}},$ (13)
+    where the DM yield computed at the standard freeze-out instant $y_{f}^{S}$ is
+    $Y_{f}^{S}=\left(1+\delta_{f}\right)\overline{Y}\left(y^{S}_{f}\right),$ (14)
     and
-    $A^{S}_{{x_{1}}\to x_{2}}=\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,m_{\rm DM}\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>_{\rm aEWSB}}{x^{2}}\,\mathrm{d}x.$ (14)
+    $A^{S}_{{y_{1}}\to y_{2}}=\sqrt{\frac{\pi}{45}}\frac{M_{\mathrm{Pl}}}{m_{\rm DM}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>_{\rm aEWSB}}{y^{2}}\,\mathrm{d}y.$ (15)
   4. (iv)
     $Y^{\infty}_{I}$ : represents the present-day yield computed in the improved approach, and is defined by
-    $Y^{\infty}_{I}=\frac{Y^{I}_{f}}{1+Y^{I}_{f}A^{I}_{{x^{I}_{f}}\to\infty}},$ (15)
-    where the DM yield computed at the improved freeze-out instant $x_{f}^{I}$ is
-    $Y_{f}^{I}=\left(1+\delta_{f}\right)\overline{Y}\left(x^{I}_{f}\right),$ (16)
+    $Y^{\infty}_{I}=\frac{Y^{I}_{f}}{1+Y^{I}_{f}A^{I}_{{y^{I}_{f}}\to\infty}},$ (16)
+    where the DM yield computed at the improved freeze-out instant $y_{f}^{I}$ is
+    $Y_{f}^{I}=\left(1+\delta_{f}\right)\overline{Y}\left(y^{I}_{f}\right),$ (17)
     and
-    $A^{I}_{{x_{1}}\to x_{2}}$ $=$ $\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,\widetilde{m}_{\mathrm{DM}}\,\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm bEWSB}}{x^{2}}\,\Theta(x_{\rm{EWSB}}-x)\,\mathrm{d}x$ (17)
-    $+\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,m_{\mathrm{DM}}\,\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm aEWSB}}{x^{2}}\,\Theta(x-x_{\rm{EWSB}})\,\mathrm{d}x.$
+    $A^{I}_{{y_{1}}\to y_{2}}$ $=$ $\sqrt{\frac{\pi}{45}}\frac{M_{\mathrm{Pl}}}{\widetilde{m}_{\rm DM}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm bEWSB}}{y^{2}}\,\Theta(y_{\rm{EWSB}}-y)\,\mathrm{d}y$ (18)
+    $+\ \sqrt{\frac{\pi}{45}}\frac{M_{\mathrm{Pl}}}{m_{\rm DM}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm aEWSB}}{y^{2}}\,\Theta(y-y_{\rm{EWSB}})\,\mathrm{d}y.$
   5. (v)
@@
 $\alpha$ $\equiv$ $\frac{1}{Y^{S}_{f}}-\frac{1}{Y^{I}_{f}},$ (22)
-$\beta$$\equiv$ $A^{I}_{{x^{I}_{f}}\to\infty}-A^{S}_{{x^{S}_{f}}\to\infty}.$ (22)
+$\beta$ $\equiv$ $A^{I}_{{y^{I}_{f}}\to\infty}-A^{S}_{{y^{S}_{f}}\to\infty}.$ (23)
 The sign of $\delta\Omega h^{2}$ is determined by the sign of $\left(\alpha-\beta\right)$.
@@
 Looking carefully at Eqs.
-(14) and (17), one may conclude that the quantity $A_{x_{\rm EWSB}\to\infty}$ is the same in both approaches.
-Further assuming both $\sigma v\right>$’s remain constant over temperature from the decoupling to EWSB, $\beta$ can be further simplified to [^3] In the illustrative model presented below, we have explicitly checked that Eq.
-(24) correctly predicts the sign of $\delta\Omega h^{2}$ for most (but not all) points in the parameter space.
-.
-$\beta$ $=A^{I}_{{x^{I}_{f}}\to x_{\rm EWSB}}-A^{S}_{{x^{S}_{f}}\to x_{\rm EWSB}}$ (23)
-[…]
+(15) and (18), one may conclude that the quantity $A_{y_{\rm EWSB}\to\infty}$ is the same in both approaches.
 Quantifying the relative error in this manner highlights the interplay between two significant factors: the annihilation cross-sections and the scalar masses, which together contribute to the shift in the final relic densities between the two approaches.
@@
 III.1 Model definition
-We extend the SM scalar sector by two real scalar singlets $\phi$ and $\chi$ [17, 30, 33]. The new scalars are stabilized by a $Z_{2}\times Z_{2}^{\prime}$ symmetry such that
+We extend the SM scalar sector by two real scalar singlets $\phi$ and $\chi$ Cline:2013gha ; Steigman:2012nb ; Basak:2021tnj .
+The new scalars are stabilized by a $Z_{2}\times Z_{2}^{\prime}$ symmetry such that
 $\chi\overset{Z_{2}}{\longrightarrow}-\chi\ \ \ \ \ \ \ \ {\rm and}\ \ \ \ \ \ \ \phi\overset{Z_{2}^{\prime}}{\longrightarrow}-\phi\,.$ (24)
@@
 $m_{h_{1}},\ m_{h_{2}},\ m_{\chi},\ v_{H},\ \sin\theta,\ \lambda_{\chi},\ \lambda_{H\phi},\ \lambda_{H\chi},\ \lambda_{\phi\chi},$ (33)
-where we fix $v_{H}=246$ GeV, and identify $h_{1}$ with the $125$ GeV scalar found at the LHC [34, 35].
+where we fix $v_{H}=246$ GeV, and identify $h_{1}$ with the $125$ GeV scalar found at the LHC ATLAS:2012yve ; CMS:2012qbp .
 Note that the values of these parameters are chosen such that the scalar potential has a minimum both before and after EWSB.
@@
 Figure 1: Annihilation channels of scalar particles before EWSB
-[…]
+If both components were to freeze out after EWSB, then, the first line of Eq.
+(18) would vanish, leading to $A^{I}_{{y_{1}}\to y_{2}}=A^{S}_{{y_{1}}\to y_{2}}$ of Eq.
+(15), and the two approaches would coincide.
+This would only be achieved if the mass hierarchies were $m_{\chi},m_{\phi}\lesssim\ 4\,{\rm TeV}$, due to the reason mentioned in Eq.
+(10).
+To simplify the analysis conservatively, in our study, we impose $m_{\phi}\lesssim\,1\ {\rm TeV}$, so that $\phi$ remains in the thermal bath before EWSB.
+As a result, $Y_{\phi}$ always traces $\overline{Y}_{\phi}$ until EWSB occurs.
+On the other hand, we require $\widetilde{m}_{\chi}\gtrsim 4\ {\rm TeV}$, such that $Y_{\chi}$ saturates before EWSB.
+Thus, the treatment of $\chi$ in the standard and improved approaches will have a noticeable physical consequence.
 When $T<T_{\mathrm{EW}}$, EWSB kicks in and the scalars $H$ and $\phi$ acquire their respective vevs, the $Z_{2}^{\prime}$ symmetry is explicitly broken, no longer rendering $\phi$ as a DM component.
@@
 As explained, the upper bound on the mass of $h_{2}$ is chosen to ensure it remains in thermal equilibrium with the SM bath until EWSB occurs.
-The range considered for the mixing angle $\theta$ is consistent with the latest experimental bounds on heavy Higgs-like scalar mixing [36].
+The range considered for the mixing angle $\theta$ is consistent with the latest experimental bounds on heavy Higgs-like scalar mixing ATLAS:2021pkb .
 Regarding quartic couplings, we scanned over their typical freeze-out values, ensuring they respect perturbativity and the boundedness from below conditions of App.
@@
 III in the electroweak broken phase.
-We used micrOMEGAs 6.0.5 [32] to numerically integrate the Boltzmann Equation and track the evolution of the yield of $\chi$ from $T=10^{5}$ GeV to $T=10^{-4}$ GeV.
+We used micrOMEGAs 6.0.5 Alguero:2023zol to numerically integrate the Boltzmann Equation and track the evolution of the yield of $\chi$ from $T=10^{5}$ GeV to $T=10^{-4}$ GeV.
 The standard relic density was then calculated using Eq.
@@
 Green points have $\Omega h^{2}|_{I}\approx 0.120$ and are allowed by Direct Detection constraints.
-[…]
+In Fig.
+3, we scan over the parameter space and plot, in gray, the relative difference ($\delta\Omega h^{2}$) between the standard and improved approaches as a function of the dark matter mass, over the range 11001–100 TeV.
+Here, the gray vertical line marks the inverse of the EWSB temperature.
+Points with relic density computed in the improved approach and lying within the $3\sigma$ range of the PLANCK measurement Planck:2018vyg are shown in red.
+As expected, $\delta\Omega h^{2}$ begins to deviate noticeably from zero as $m_{\chi}$ approaches 4 TeV — the approximate lower bound on the dark matter mass required for it to decouple from the SM bath before electroweak symmetry breaking, as discussed in Sec.
+II.
+A subset of the red points, marked with green, satisfy the additional requirement that the direct detection (DD) cross-section lies below the current experimental limits Bhoonah:2020fys ; LZ:2024psa .
+We provide an explicit formula for the spin-independent scattering cross-section in App.
+C.
+For $m_{\chi}\lesssim 10\ {\rm TeV}$, we applied the latest DD exclusion limits on the spin-independent scattering cross-section off nuclei from the 2024 LUX-ZEPLIN (LZ) experiment LZ:2024zvo .
+We found that the green points tend to favour small values of $\delta\Omega h^{2}$, reflecting the fact that in this simplified model, both the relic density and the DD cross-section are controlled by the scalar quartic couplings $\lambda_{H\chi}$ or $\lambda_{\phi\chi}$.
+These couplings need to be large to achieve the observed relic density, but small enough to pass DD constraints.
+However, we expect these exclusions to be less restrictive in more complete models, where the relic abundance and DD channels are not necessarily correlated.
+In Table.
+1, we show two red benchmark points with $m_{\chi}\lesssim\ 10\ {\rm TeV}$ with non-zero $\delta\Omega h^{2}$ that have DD cross-sections below the 2024 LZ bounds LZ:2024zvo .
 Masses Mixing and Couplings Relic Abundance
@@
 5.737 984 0.001 0.003 0.003 3.635 0.122721 0.125122 1.95
-8.281 467 0.036 $-0.095$ $-0.248$ 5.103 0.118672 0.119259 0.49
-[…]
+8.281 467 0.036 0.095-0.095 0.248-0.248 5.103 0.118672 0.119259 0.49
+Table 1: Benchmark points with $m_{\chi}\lesssim\,10\ {\rm TeV}$ with finite $\delta\Omega h^{2}$ and direct detection cross-sections below LZ limits.
 Notice that the sign of $\delta\Omega h^{2}$ may either be positive or negative, as a consequence of the improved relic density lying above or below the standard relic density, respectively.
@@
 For each value of $\lambda_{H\chi}$, we trace in the bottom row of plots the corresponding temperature evolution of $\langle\sigma v\rangle_{I}$ with a black solid line, and that of $\langle\sigma v\rangle_{S}$ with a red dashed line.
-It is interesting to note that $\chi$ always decouples from the SM bath earlier than in the improved approach, implying the relation $x_{f}^{S}\lesssim x_{f}^{I}$.
+It is interesting to note that $\chi$ always decouples from the SM bath earlier than in the improved approach, implying the relation $y_{f}^{S}\lesssim y_{f}^{I}$.
 Yet, regardless of the value of $\lambda_{H\chi}$, we always observe $\chi$ decoupling from the SM bath when $T>T_{\rm EWSB}$, as expected from the DM mass requirement in Eq.
@@
 At freeze-out, the standard and improved yields satisfy $Y_{f}^{S}>Y_{f}^{I}$, which implies a negative value for $\alpha$.
-On the other hand, before EWSB, we notice that $\langle\sigma v\rangle_{S}\lesssim\langle\sigma v\rangle_{I}$, which implies that $A^{S}_{{x^{I}_{f}}\to\infty}<A^{I}_{{x^{I}_{f}}\to\infty}$ and results in a positive $\beta$.
+On the other hand, before EWSB, we notice that $\langle\sigma v\rangle_{S}\lesssim\langle\sigma v\rangle_{I}$, which implies that $A^{S}_{{y^{I}_{f}}\to\infty}<A^{I}_{{y^{I}_{f}}\to\infty}$ and results in a positive $\beta$.
 For clarity, the particular values of $\alpha$ and $\beta$ for this benchmark point are marked by the white circular symbol in Fig.
@@
 The ongoing and planned direct and indirect detection experiments (see Ref.
-[1] for a comprehensive review) will most certainly give rise to an increase (or even exclusion) in the upper bound of the DM mass in this simple scenario with two singlets and two independent symmetries.
+Cirelli:2024ssz for a comprehensive review) will most certainly give rise to an increase (or even exclusion) in the upper bound of the DM mass in this simple scenario with two singlets and two independent symmetries.
 However, it is easy to extend the model with, for example, an extra singlet, such that the couplings responsible for direct and indirect detection are not the same as the ones that determine the value of the relic density.
 We defer a detailed investigation of such extensions to future work.
-The study also builds a strong case for testing TeV-scale phenomena in upcoming high-energy particle colliders such as the FCC-hh [40].
+The study also builds a strong case for testing TeV-scale phenomena in upcoming high-energy particle colliders such as the FCC-hh Benedikt:2022kan .
 In fact, although models for which our approach is important have very heavy DM candidates, they could in principle be probed at future colliders.
@@
 Appendix B Boundedness from below
-For the scalar potential studied in this paper to be bounded from below, we must ensure that [41, 42]:
+For the scalar potential studied in this paper to be bounded from below, we must ensure that Kannike:2012pe ; Habibolahi:2022rcd :
 $\lambda_{H}>0\,,\quad\quad\lambda_{\phi}>0\,,\quad\quad\lambda_{\chi}>0\,.$ (50)
@@
 Appendix C Direct Detection
-The DM-nucleon spin-independent elastic scattering cross-section ($\sigma_{\mathrm{SI}}$), in this model reads [43]
+The DM-nucleon spin-independent elastic scattering cross-section ($\sigma_{\mathrm{SI}}$), in this model reads Cerdeno:2010jj
 $\sigma_{\mathrm{SI}}$ $=$ $\frac{1}{\pi}\left(\frac{m_{N}^{2}}{m_{N}+m_{\chi}}\right)^{2}\frac{f_{N}^{2}}{v_{H}^{2}}\left|\frac{\cos\theta\lambda_{h_{1}\chi\chi}}{m_{h_{1}}^{2}}-\frac{\sin\theta\lambda_{h_{2}\chi\chi}}{m_{h_{2}}^{2}}\right|^{2},$ (53)
-where $f_{N}\approx 0.308$ is the Higgs-nucleon effective coupling [44], $m_{N}\approx 0.938$ GeV is the nucleon mass, and
+where $f_{N}\approx 0.308$ is the Higgs-nucleon effective coupling Hoferichter:2017olk , $m_{N}\approx 0.938$ GeV is the nucleon mass, and
 $\lambda_{h_{1}\chi\chi}$ $=$ $\cos\theta\,v_{H}\lambda_{H\chi}+\sin\theta\,v_{\phi}\lambda_{\phi\chi},$ (54)